1)Divide the sum of -8/7 and 5/14 by their product?
2) What should be added to the sum of 3/4 and -2/-5 to get -8/15?

Answer by rfer(16322)

sum
-16/14+5/14=-11/14
product
-8/7*5/14=-40/98=-20/49
------------
3/4+2/5+x=-8/15
15/20+8/20+x=-8/15
23/20+x=-8/15
x=-8/15-23/20
x=-32/60-69/60
x=-101/60=
-------------
3/4+2/5-101/60=-8/15
45/60+24/60-101/60=-8/15